OpenStudy (naga447):

Sequences help??? Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? f(n) = 28(0.5)^n f(n) = 56(0.5)^n-1 f(n) = 56(0.5)^n f(n) = 112(0.5)^n-1

OpenStudy (kirbykirby):

By inspection, I notice that b), will give 56 for n=1, since \(f(1)=56(0.5)^{1-1}=56(0.5)^0=56(1)=56\) Then if you want to be sure, you can test whether n=2, gives 28, etc.

OpenStudy (naga447):

Thank you! sorry for the late reply